Evaluate ∬R(x-y)^{2}sin^{2}(x+y)d xdy, where R is the rhombus withsuccessive vertices at (π,0),(2π,π),(π,2t) and (0,π)


1
Expert's answer
2021-08-30T01:55:06-0400

"∬_R(x-y)^{2}sin^{2}(x+y)d xdy\\
∬_R((x-y)sin(x+y))^2d xdy\\
G(u,v)= (\frac{u+v}{2}, \frac{u-v}{2})\\
x= \frac{u+v}{2}, y= \frac{u-v}{2}\\
dxdy= |J(u,v)|dudv\\
J(u,v)=\begin{vmatrix}
 \frac{\partial x}{\partial u} & \frac{\partial x}{\partial v} \\
 \frac{\partial y}{\partial u} & \frac{\partial y}{\partial v}
\end{vmatrix} = \begin{vmatrix}
 \frac{1}{2} & \frac{1}{2} \\
 \frac{1}{2} & -\frac{1}{2}
\end{vmatrix} = \frac{1}{2}\\
∬_R((x-y)sin(x+y))^2d xdy= \int_0^{2 \pi}\int_0^{2 \pi}(v-\sin u) J(u,v)d udv\\
= \int_0^{2 \pi}\int_0^{2 \pi}v^2\sin^2 u (\frac{1}{2})d udv\\
= \frac{1}{2}\int_0^{2 \pi}dv\int_0^{2 \pi}\sin^2 udu\\
= \frac{1}{2}(2 \pi)\int_0^{2 \pi}(\frac{1-\cos 2 u}{2})du\\
= \frac{1}{2}(2 \pi)[\frac{1}{2}u -\frac{\sin 2 u}{u}]_0^{2 \pi}\\
= \frac{1}{2}(2\pi)( \frac{1}{2} (2 \pi))\\
=\pi^2"